# Making AI Less “Thirsty”: Uncovering and Addressing the Secret Water Footprint of AI Models

### v4

@@ 1 Introduction
$\bullet$ “Water is a finite resource, and every drop matters.” — Facebook (now Meta) Sustainability Report 2020 [1]. $\bullet$ “Fresh, clean water is one of the most precious resources on Earth … Now we’re taking urgent action to support water security and healthy ecosystems.” — Google’s Water Commitment 2023 [2]. $\bullet$ “Water is a human right and the common development denominator to shape a better future. But water is in deep trouble.” — U.N. Secretary-General António Guterres at the U.N. Water Conference 2023 [3]. $\bullet$ “Historic droughts threaten our supply of water … As the source of both life and livelihoods, water security is central to human and national security.” — The U.S. White House Action Plan on Global Water Security 2022 [4].
@@
As acknowledged in Google’s sustainability report [8] and the recent U.S. Department of Energy study [5], the expansion of AI products and services is a key driver of the rapid increase in data center water consumption. Even excluding the water usage in leased third-party colocation facilities, Google’s self-owned data centers alone directly withdrew 29 billion liters and consumed (i.e., evaporated) more than 23 billion liters of freshwater for on-site cooling in 2023, nearly 80% of which was potable water [8]. [^2] The detailed difference between water withdrawal and water consumption is presented in Section 2.1. This amount of annual water consumption even rivals that of a major household-name beverage company [9]. Importantly, Google’s data center water consumption increased by $\sim$20from 2021 to 2022 and by $\sim$17from 2022 to 2023 [8], and Microsoft saw $\sim$34and $\sim$22increases over the same periods, respectively [10]. Furthermore, according to the U.S. Department of Energy, the total annual on-site water consumption by U.S. data centers in 2028 could double or even quadruple the 2023 level, reaching approximately 150 – 280 billion liters and further stressing the water infrastructures [5].
AI represents the fastest expanding workloads in data centers [8, 5]. For example, a recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 [11], whereas a more aggressive projection by the U.S. Department of Energy predicts that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [5]. Even considering the lower estimate, the combined scope-1 and scope-2 water withdrawal of global AI is projected to reach 4.2 – 6.6 billion cubic meters in 2027, which is more than the total annual water withdrawal of 4 – 6 Denmark or half of the United Kingdom. [^3] The scope definition of water usage [12] is in line with that of carbon emissions and is discussed in Section 2.2. Our scope-2 water withdrawal (and consumption when applicable) is for location-based electricity generation throughout the paper. Large data centers often adopt sustainability programs (e.g., renewable purchasing agreements) to offset their location-based electricity usage and thus may have lower market-based carbon and water footprints. Simultaneously, a total of 0.38 – 0.60 billion cubic meters of water will be evaporated and considered “consumption” due to the global AI demand in 2027. Moreover, these global estimates will be exceeded by the total water withdrawal and consumption attributed to AI in the U.S. alone in 2028 if the U.S. Department of Energy’s projection comes to fruition.
@@
To ensure that the growth in AI does not exacerbate the global water stresses or outweigh the environmental benefits it provides, it is critical to uncover and address AI’s hidden water footprint amid the increasingly severe freshwater scarcity crisis, worsened extended droughts, and quickly aging public water infrastructure. The urgency can also be reflected in part by the recent commitment to “Water Positive by 2030” from industry leaders, including Google [8] and Microsoft [10], by policy guidelines and legislative efforts to mitigate AI’s water consumption [14, 15], and by the inclusion of water consumption as a key environmental metric into the first international standard on sustainable AI to be published by the ISO/IEC [16]. Moreover, setting targets for minimizing water consumption is included as part of the recent U.S. executive order for advancing AI data center infrastructures [17].
In this paper, we advocate for a holistic approach to sustainable AI that extends beyond the carbon footprint to also address the water footprint. Specifically, we present a principled methodology to estimate AI’s total water footprint, including both operational water and embodied water. By taking the GPT-3 model with 175 billion parameters as an example [18], we show that training GPT-3 in Microsoft’s U.S. data centers can consume a total of 5.4 million liters of water, including 700,000 liters of scope-1 on-site water consumption. Additionally, GPT-3 needs to “drink” (i.e., consume) a 500ml bottle of water for roughly 10 – 50 medium-length responses, depending on when and where it is deployed. Note that our estimate of inference water consumption for GPT-3 is conservative, and the actual water consumption could be several times higher.
Next, we emphasize the need for increasing transparency of AI’ water footprint, including disclosing more information about operational data and keeping users informed of the runtime water efficiency. We show that WUE (Water Usage Effectiveness, a measure of water efficiency) varies both spatially and temporally, suggesting that “when” and “where” to perform AI training can significantly affect the water footprint. Finally, we highlight the necessity of holistically addressing the water footprint along with the carbon footprint to enable truly sustainable AI — the water footprint of AI can no longer stay under the radar.
@@ 2.2.2 Scope-2 Water Usage
In many countries, thermoelectric power is among the top sectors in terms of water withdrawal and water consumption [12]. Thus, similarly to scope-2 carbon emissions, data centers are accountable for off-site scope-2 water usage associated with electricity consumption, which forms part of the “true water cost of data centers,” as highlighted by the U.S. Department of Energy [5].
@@ 3.2 Embodied Water Footprint
By adding up the operational and embodied water footprints, we obtain the total water footprint as $WaterTotal=\sum_{t=1}^{T}e_{t}\cdot\left[\rho_{s1,t}+\theta_{t}\cdot\rho_{s2,t}\right]+\frac{T\cdot W}{T_{0}}$. In practice, to obtain a rough estimate, we can use the average values for the annualized WUE and the estimated AI server energy consumption.
@@ 3.3.1 Training
GPT-3 was trained and deployed by OpenAI in Microsoft’s data centers, with an estimated training energy of 1287 MWh [36]. In line with the practice of estimating the carbon footprint, we use the most recent annualized average on-site PUE and WUE for each location, as reported by Microsoft [34, 35]. For power plant water efficiency, different references may provide different estimates of EWIF. Thus, for consistency across regions, we use the EWIF provided by [12] to estimate scope-2 water consumption, as it employs the same methodology for calculating EWIF across a large number of regions. Moreover, a large number of Microsoft’s data centers are located in the U.S., where the average EWIF provided by [12] is 3.14 L/kWh and significantly lower than 4.35 L/kWh reported by the recent U.S. Department of Energy study [5]. The specific location for training GPT-3 is not public. Thus, we consider Microsoft’s different data center locations, while excluding Singapore and Taiwan since the EWIF data for these regions is not available in [12].
@@ 3.3.2 Inference
As a representative usage scenario for an LLM, we consider a conversation task, which typically includes a CPU-intensive prompt phase that processes the user’s input (a.k.a., prompt) and a memory-intensive token phase that produces outputs [37]. More specifically, we consider a medium-sized request, each with approximately $\leq$800 words of input and 150 – 300 words of output [37]. The official estimate shows that GPT-3 consumes an order of 0.4 kWh electricity to generate 100 pages of content (e.g., roughly 0.004 kWh per page) [18]. Thus, we consider 0.004 kWh as the per-request server energy consumption for our conversation task. The PUE, WUE, and EWIF are the same as those used for estimating the training water consumption.
Our estimate of inference water consumption for GPT-3 is on the conservative side, and the actual water consumption could be several times higher. Concretely, when service level objectives (SLOs) for LLM response time are considered in real enterprise-grade Nvidia DGX H100 systems for conversation tasks, the inference server energy consumption for a much smaller model (e.g., Llama-3-70B) to process each medium-sized request is already $\sim$0.010 kWh when using a state-of-the-practice LLM inference solution and accounting for the non-GPU server overhead [37]. When considering the Falcon-180B model with a comparable size to GPT-3-175B, the server energy consumption will even reach $\sim$0.016 kWh for processing a medium-sized request [37]. Additionally, our EWIF for the U.S. (i.e., 3.14L/kWh on average) is conservatively chosen and significantly lower than 4.35L/kWh recently reported by [5].
While there is no official information on its resource consumption, the subsequent GPT-4 model is expected to consume substantially more energy and water than GPT-3 for processing the same request [38]. Nonetheless, we emphasize that the on-site WUE of Microsoft’s data centers is already among the lowest in the industry. With continued efforts to reduce AI’s computational demand and improve the overall water efficiency, the water consumption per request may decrease in the future, while the total water consumption is likely to rise further as a result of the growing demand for AI services and the increasing scale of AI applications [5].
@@ 4.1 More Transparency and Comprehensive Reporting
Moreover, a comprehensive understanding and reporting of AI’s scope-2 water consumption associated with electricity generation remain limited. Although data centers have increasingly adopted climate-conscious cooling system designs to minimize on-site water consumption [8, 26, 24], these efforts primarily focus on scope-1 water usage while largely overlooking scope-2 impacts. Just as addressing scope-2 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 water consumption to reduce AI’s “true water cost”, as noted by the U.S. Department of Energy study [5]. To better reflect the true impacts of data centers on water resources, some technology companies such as Meta have begun to include scope-2 water consumption in their sustainability reports [25]. We recommend the reporting of scope-2 water consumption as a standard practice. This approach makes the off-site water consumption visible to AI model developers as well as end users and can unlock new opportunities for demand-side flexibility, thereby alleviating the overall strain on water resources.
@@ 5 Conclusion
A recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 based on the GPU shipment [11], whereas a more aggressive projection by the U.S. Department of Energy predicts that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [5]. Based on the former and more conservative projection, we estimate the potential water usage for global AI in 2027, while noting that our global estimates will be exceeded by the water usage attributed to AI in the U.S. alone in 2028 if the U.S. Department of Energy’s projection comes to fruition.
Scope-1 water usage. The scope-1 water efficiency depends on a variety of factors, including the cooling system designs, climate conditions, and operational settings. To set the global scope-1 water efficiency, we utilize the annualized water efficiencies reported by two leading data center operators, Google and Equinix, in their latest sustainability reports [8, 40]. Specifically, for on-site scope-1 water withdrawal, we assume 1.2 L/kWh, which results in a total scope-1 water withdrawal of 0.11 – 0.16 billion cubic meters. Similarly, assuming 1.0 L/kWh for the global scope-1 water consumption efficiency, we obtain a total on-site scope-1 water consumption of 0.09 – 0.14 billion cubic meters. Note that Google and Equinix both operate data centers globally, but represent two distinct categories of data centers: hyperscale data centers (Google) and multi-tenant colocation data centers (Equinix). According to the U.S. Department of Energy [5], these two types of data centers collectively account for the vast majority of data center energy consumption in the U.S., with colocation data centers consuming slightly more energy than hyperscalers.
Scope-2 water usage. As noted by the U.S. Department of Energy [5], scope-2 water usage is part of the true water cost of data centers. The U.S. average electricity water withdrawal and consumption intensity factors are both lower than the global averages [12]. Thus, in our estimate, we use the U.S. average electricity water withdrawal intensity factor 43.83 L/kWh [41], and electricity water consumption intensity factor 3.14 L/kWh [12], respectively. Note that, since [12] includes hydropower in the calculation, it has a higher electricity water withdrawal factor than the U.S. Energy Information Administration’s calculation (i.e., 386.07 L/kWh vs. 43.83 L/kWh for the U.S.). Moreover, our value of 3.14 L/kWh for the U.S. average water consumption factor is lower than 4.35 L/kWh reported by the U.S. Department of Energy [5], as well as lower than Meta’s global electricity water consumption intensity factor of 3.70 L/kWh in 2024 (i.e., 55,475 megaliters divided by 14,975,435 MWh) [25]. Therefore, our choices of 43.83 L/kWh and 3.14 L/kWh for electricity water withdrawal and consumption intensity factors are both on the conservative side, which can partly absorb potential over-estimates of global AI’s energy demand in 2027 provided by [11].
@@
The estimates of global AI’s water usage in 2027 are naturally subject to uncertainties, e.g., the future water efficiency may differ from the current value we use. Nonetheless, we emphasize that our estimates are on the conservative side. For example, considering a higher estimate, the U.S. Department of Energy’s scope-1 water consumption attributed to AI in the U.S. alone could exceed 0.2 billion cubic meters in 2028 [5]. Moreover, based on the reported scope-2 water consumption efficiency, the combined scope-1 and scope-2 water consumption attributed to AI in the U.S. alone is projected to reach up to about 2 billion cubic meters in 2028 [5], which is significantly higher than our estimate of global AI’s total water consumption in 2027.

### v5

@@ 1 Introduction
As acknowledged in Google’s sustainability report [4] and the recent U.S. data center energy report [1], the expansion of AI products and services is a key driver of the rapid increase in data center water consumption. Even excluding the water usage in leased third-party colocation facilities, one technology company’s self-owned data centers alone directly withdrew 29 billion liters and consumed (i.e., evaporated) more than 23 billion liters of freshwater for on-site cooling in 2023, nearly 80% of which was potable water [4]. [^2] The detailed difference between water withdrawal and water consumption is presented in Section 2.1. This amount of annual water consumption even rivals that of a major household-name beverage company [5]. Importantly, the company’s data center water consumption increased by $\sim$20from 2021 to 2022 and by $\sim$17from 2022 to 2023 [4], and another technology company’s data center water consumption saw $\sim$34and $\sim$22increases over the same periods, respectively [6]. Furthermore, according to the recent U.S. data center energy report, the total annual on-site water consumption by U.S. data centers in 2028 could double or even quadruple the 2023 level, reaching approximately 150 – 280 billion liters and further stressing the water infrastructures [1].
AI represents the fastest expanding workloads in data centers [4, 1]. For example, a recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 [7], whereas a more aggressive projection by the recent U.S. data center energy report predicts that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [1]. Even considering the lower estimate, the combined scope-1 and scope-2 water withdrawal of global AI is projected to reach 4.2 – 6.6 billion cubic meters in 2027, which is more than the total annual water withdrawal of 4 – 6 Denmark or half of the United Kingdom. [^3] The scope definition of water usage [8] is in line with that of carbon emissions and is discussed in Section 2.2. Our scope-2 water withdrawal (and consumption when applicable) is for location-based electricity generation throughout the paper. Large data centers often adopt sustainability programs (e.g., renewable purchasing agreements) to offset their location-based electricity usage and thus may have lower market-based carbon and water footprints. Simultaneously, a total of 0.38 – 0.60 billion cubic meters of water will be evaporated and considered “consumption” due to the global AI demand in 2027. Moreover, these global estimates will be exceeded by the total water withdrawal and consumption attributed to AI in the U.S. alone in 2028 if the projection in [1] comes to fruition.
@@
To ensure that the growth in AI does not exacerbate the global water stresses or outweigh the environmental benefits it provides, it is a critical time to uncover and address AI’s hidden water footprint amid the increasingly severe freshwater scarcity crisis, worsened extended droughts, and quickly aging public water infrastructure. The urgency can also be reflected in part by the recent commitment to “Water Positive by 2030” from industry leaders, including Google [4] and Microsoft [6], and by the inclusion of water footprint as a key metric into the world’s first international standard on sustainable AI to be published by the ISO/IEC [10].
In this paper, we advocate for a holistic approach to sustainable AI that extends beyond the carbon footprint to also address the water footprint. Specifically, we present a principled methodology to estimate AI’s total water footprint, including both operational water and embodied water. By taking the GPT-3 model with 175 billion parameters as an example [11], we show that training GPT-3 in Microsoft’s U.S. data centers can consume a total of 5.4 million liters of water, including 700,000 liters of scope-1 on-site water consumption. Additionally, GPT-3 needs to “drink” (i.e., consume) a 500ml bottle of water for roughly 10 – 50 medium-length responses, depending on when and where it is deployed.
Next, we show that WUE (Water Usage Effectiveness, a measure of water efficiency) varies both spatially and temporally, suggesting that judiciously deciding “when” and “where” to train a large AI model can significantly cut the water footprint. We also emphasize the need for increasing transparency of AI’ water footprint, including disclosing more information about operational data and keeping users informed of the runtime water efficiency. Finally, we highlight the necessity of holistically addressing water footprint along with carbon footprint to enable truly sustainable AI — the water footprint of AI can no longer stay under the radar.
@@ 2.2.2 Scope-2 Water Usage
In many countries, thermoelectric power is among the top sectors in terms of water withdrawal and water consumption [8]. Thus, similarly to scope-2 carbon emissions, data centers are accountable for off-site scope-2 water usage associated with electricity consumption, which forms part of the “true water cost of data centers,” as highlighted by the recent U.S. data center energy report [1].
@@ 3.2 Embodied Water Footprint
By adding up the operational and embodied water footprints, we can obtain the total water footprint as $WaterTotal=\sum_{t=1}^{T}e_{t}\cdot\left[\rho_{s1,t}+\theta_{t}\cdot\rho_{s2,t}\right]+\frac{T\cdot W}{T_{0}}$. In practice, to obtain a rough estimate, we can use the average values for the annualized WUE and the estimated AI server energy consumption.
@@ 3.3.1 Training
GPT-3 was trained and deployed by OpenAI in Microsoft’s data centers, with an estimated training energy of 1287 MWh [29]. In line with the practice of estimating the carbon footprint, we use the most recent annualized average on-site PUE and WUE for each location, as reported by Microsoft [27, 28]. For power plant water efficiency, different references may provide different estimates of EWIF. Thus, for consistency across regions, we use the EWIF provided by [8] to estimate scope-2 water consumption, as it employs the same methodology for calculating EWIF. Moreover, a large number of Microsoft’s data centers are located in the U.S., where the average EWIF provided by [8] is 3.14 L/kWh and significantly lower than 4.35 L/kWh reported by the recent study [1]. The specific location for training GPT-3 is not public. Thus, we consider Microsoft’s different data center locations, while excluding Singapore and Taiwan as EWIF data for these regions is not available in [8].
@@ 3.3.2 Inference
As a representative usage scenario for an LLM, we consider a conversation task, which typically includes a CPU-intensive prompt phase that processes the user’s input (a.k.a., prompt) and a memory-intensive token phase that produces outputs [30]. More specifically, we consider a medium-sized request, each with approximately $\leq$800 words of input and 150 – 300 words of output [30]. The official estimate indicates that GPT-3 consumes an order of 0.4 kWh of electricity to generate 100 pages of content, equivalent to roughly 0.004 kWh per page [11]. While no details are provided, the estimate likely considers only the GPU energy used during token generation.
To account for both the prompt phase and the non-GPU energy consumption of servers, we assume a per-request server energy consumption of 0.004 kWh for our conversation task. The PUE, WUE, and EWIF are the same as those used for estimating the training water consumption. Our estimate of inference water consumption for GPT-3 is on the conservative side, and the actual water consumption could be several times higher. Specifically, when considering service level objectives (SLOs) for LLM response times in enterprise-grade Nvidia DGX H100 systems for conversation tasks, the inference server energy consumption for a much smaller model (e.g., Llama-3-70B) is already approximately 0.010 kWh per medium-sized request when using a state-of-the-art LLM inference solution and accounting for non-GPU server overhead [30]. For the Falcon-180B model, which is comparable in size to GPT-3-175B, the server energy consumption reaches approximately 0.016 kWh per medium-sized request [30]. Furthermore, we emphasize that Microsoft’s data centers already have some of the lowest on-site WUE in the industry. If the same model is deployed in a third-party colocation data center, the scope-1 direct water consumption is expected to be several times higher. Additionally, our EWIF for the U.S. (3.14 L/kWh) is conservative and significantly lower than the 4.35 L/kWh recently reported by [1].
While no official information is available on the resource consumption, some subsequent models like GPT-4 could consume substantially more energy and water than GPT-3 for processing the same request [31, 32]. With continued efforts to reduce AI’s computational demand and improve the overall water efficiency, the water consumption per request may decrease in the future. However, the total water consumption is likely to continue rising due to the growing demand for AI services and the increasing scale of AI applications [1].
@@ 4.1 More Transparency and Comprehensive Reporting
Moreover, a comprehensive understanding and reporting of AI’s scope-2 water consumption associated with electricity generation remain limited. Although data centers have increasingly adopted climate-conscious cooling system designs to minimize on-site water consumption [4, 19, 17], these efforts primarily focus on scope-1 water usage while largely overlooking scope-2 impacts. Just as addressing scope-2 carbon emissions is important for mitigating climate change, it is equally crucial to address scope-2 water consumption to reduce AI’s “true water cost”, as noted by the recent U.S. data center energy report [1]. To better reflect the true impacts of data centers on water resources, some technology companies such as Meta have begun to include scope-2 water consumption in their sustainability reports [18]. We recommend the reporting of scope-2 water consumption as a standard practice. This approach makes the off-site water consumption visible to AI model developers as well as end users and can unlock new opportunities for demand-side flexibility, thereby alleviating the overall strain on water resources.
@@ 5 Conclusion
A recent study suggests that the global AI could consume 85 – 134 TWh of electricity in 2027 based on the GPU shipment [7], whereas a more aggressive estimate by the U.S. data center energy report projects that AI servers’ electricity consumption in the U.S. alone will surpass 150 – 300 TWh in 2028 [1]. Based on the former and more conservative projection, we estimate the potential water usage for global AI in 2027, while noting that our global estimates will be exceeded by the water usage attributed to AI in the U.S. alone in 2028 if the projection in [1] comes to fruition.
Scope-1 water usage. The scope-1 water efficiency depends on a variety of factors, including the cooling system designs, climate conditions, and operational settings. To set the global scope-1 water efficiency, we utilize the annualized water efficiencies reported by two leading data center operators, Google and Equinix, in their latest sustainability reports [4, 34]. Specifically, for on-site scope-1 water withdrawal, we assume 1.2 L/kWh, which results in a total scope-1 water withdrawal of 0.11 – 0.16 billion cubic meters. Similarly, assuming 1.0 L/kWh for global scope-1 water consumption efficiency, we obtain a total on-site scope-1 water consumption of 0.09 – 0.14 billion cubic meters. Note that Google and Equinix both operate data centers globally, but represent two distinct categories of data centers: hyperscale data centers (Google) and multi-tenant colocation data centers (Equinix). According to the recent U.S. data center energy report [1], these two types of data centers collectively account for the vast majority of data center energy consumption in the U.S., with colocation data centers consuming slightly more energy than hyperscalers.
Scope-2 water usage. As noted by the recent U.S. data center energy report [1], scope-2 water usage is part of the true water cost of data centers. The U.S. average electricity water withdrawal and consumption intensity factors are both lower than the global averages [8]. Thus, in our estimate, we use the U.S. average electricity water withdrawal intensity factor 43.83 L/kWh [35], and electricity water consumption intensity factor 3.14 L/kWh [8], respectively. Note that, since [8] includes hydropower in the calculation, it has a higher electricity water withdrawal factor than the U.S. Energy Information Administration’s calculation (i.e., 386.07 L/kWh vs. 43.83 L/kWh for the U.S.). Moreover, our value of 3.14 L/kWh for the U.S. average water consumption factor is lower than 4.35 L/kWh reported by [1], as well as lower than Meta’s global electricity water consumption intensity factor of 3.70 L/kWh in 2024 (i.e., 55,475 megaliters divided by 14,975,435 MWh) [18]. Therefore, our choices of 43.83 L/kWh and 3.14 L/kWh for electricity water withdrawal and consumption intensity factors are both on the conservative side, which can partly absorb potential over-estimates of global AI’s energy demand in 2027 provided by [7].
@@
The estimates of global AI’s water usage in 2027 are naturally subject to uncertainties, e.g., the future water efficiency may differ from the current value we use. Nonetheless, we emphasize that our estimates are on the conservative side. For example, according to the U.S. data center energy report, the scope-1 water consumption attributed to AI in the U.S. alone could exceed 0.2 billion cubic meters in 2028 [1]. Moreover, based on the reported scope-2 water consumption efficiency, the combined scope-1 and scope-2 water consumption attributed to AI in the U.S. alone is projected to reach up to about 2 billion cubic meters in 2028 [1], which is significantly higher than our estimate of global AI’s total water consumption in 2027.